Use the financial statements of Heifer Sports Inc. to find the information below for Heifer. (Use 365 days a year. Round all answers to 2 decimal places except $ amounts.) Income Statement Sales Cost of goods sold Depreciation Selling & admin. expenses EBIT Interest expense Taxable income Taxes Net income 2017 $ 5,650,000 2,950,000 291,206 1,560,000 $848,800 146,000 $702,800 315,400 $ 387,400 Balance Sheet, Year-End AssetS 2017 2016 Cash Accounts receivable Inventory $ 47,300 640,000e 475,100 60,000 1,037, 500 721,700 Total current assets Fixed assets $ 1,162,4001.819 - 3,024,000 $ 4,186,400 4,231,000 $ 6,050,200 Total assets Liabilities and Stockholders Equity Accounts payable Short-term debt $ 332,500 493,000 $ 705,000 863,300 Total current liabilities Long-term bonds 825,500568 3- 2,205,1003,453,200_ $ 3,030,600 Total liabilities Common stock Retained earnings $ 5,021,800 $ 313,800 714,600 $ 1,028,400 $6,050, 200 $ 313,800 842,00e Total stockholders equity $1,155,801 Total liabilities and stockholders equity $ 4,186,400 Answer is not complete Inventory turnover ratio in 2017 b. Debt/equity ratio in 2017 c. Cash flow from operating activities in 2017 4.93 2.62 S950,200Balance Sheet, Year-End Assets 2017 2016 Cash Accounts receivable Inventory $47,300 640,000e 475,100 $ 1,162,400 $ 60,000 1,037,500 721,700 23 7100_ Total current assets Fixed assets $ 1,819,200 3,024,0004231,900 Total assets $ 4,186,400 $6,050,200 Liabilities and Stockholders Equity Accounts payable Short-term debt $ 332,500 493,000 $ 825,500 $ 705,000 863,300 $ 1,568,300 2,205,1003,453-300 $ 5,021,800 313,880313,888 714,600 $ 1,028,400 $6,050,200 Total current liabilities Long-term bonds Total liabilities Common stock Retained earnings $ 3,030,600 Total stockholders equity $1,155,800 $ 4,186,400 Total liabilities and stockholders equity Answer is not complete Inventory turnover ratio in 2017 b.Debt/equity ratio in 2017 c. Cash flow from operating activities in 2017 4.93 2.62 S950,200 54.18days Average collection period Asset turnover ratio Interest coverage ratio 1.10 5.81 g. Operating profit margin h. Return on equity 15.0 % 33.51 % Compound leverage ratio 3.88 k.Net cash provided by operating activities
Answer #1

(a) The answer for operating margin is wrong because in this context, operting income of the firm includes depreciation. Hence, Operating Income = EBITDA =  (EBIT + Depreciation)

Operating Income = 848800 + 291200 = $ 1140000 and Revenue = $ 5650000

Therefore, Operating Margin = (1140000 / 5650000) x 100 ~ 20.1769 % ~ 20.18 %

(b) Return on Equity (at the end of Time T) = Net Income (at the end of Time T) / Shareholder's Equity (at the end of Time T-1) (because ROE measures the return (net income) generated for a given amount of equity already invested).

Therefore, ROE = Net Income 2017 / Shareholder's Equity 2016 = 387400 / 1028400 ~ 0.3767 or 37.67 %